# UTrace: Poisoning Forensics for Private Collaborative Learning

### v1

@@ A.2 True Unlearning Failure Modes
Given the deficiencies of approximate unlearning demonstrated in Section A.1, it may seem tempting to replace it with an unlearning mechanism which more accurately approximates leave-one-out retraining. However, we show that in fact true unlearning exacerbates the ability of the adversary to hide poisons among multiple poisoned parties.
@@
In the case when the auditor produces the unlearned models $\theta_{i}^{\prime}$ by retraining from scratch for each user $i$, each time leaving out the user dataset $D_{i}$, the effect of duplicated poisons across multiple user datasets is exacerbated. We consider a minimal theoretical example demonstrating this possibility. Consider two distributions $\mu_{c}$ and $\mu_{p}$ over $\mathcal{X}\times\mathcal{Y}$, representing the clean and poisoned distributions, respectively. If the distributions $\mu_{c}$ and $\mu_{p}$ are compatible with each other, and if the model space $\Theta$ is sufficiently rich to simultaneously learn both distributions, then learning on any nontrivial mixture of those distributions yields the same set of loss-minimizing classifiers. Intuition for this idea is given Figure 5. We formalize this idea below:
@@
Implications of Theorem A.2. The distributions $\mu_{c}$ and $\mu_{p}$ in Theorem A.2 can be replaced with the empirical distributions $D_{i}$ for each user. Interpreted this way, the theorem describes a simplified setting where each user dataset is one of two possible datasets, $D_{c}$ or $D_{p}$. If the necessary assumptions are satisfied, performing true unlearning on any individual dataset that is duplicated across multiple users corresponds to setting a different $\alpha\in(0,1)$ (namely, $\alpha=\frac{a-1}{a+b-1}$, where $a$ and $b$ are the multiplicities of each dataset and we have removed one from those counted by $a$).
@@ B.1 Ideal Functionalities
Functionality F.1 $\mathcal{F}_{\mathsf{fp-mul}}$ Input: The functionality receives inputs $[\![\,a\,]\!]$ and $[\![\,b\,]\!]$. Output: Compute the following 1. Reconstruct $a$ and $b$ to compute $ab^{\prime}\leftarrow a\cdot b$. 2. Truncate the lower $f$ bits of $ab^{\prime}$ to get $ab$. 3. Distribute shares $[\![\,ab\,]\!]$ to the parties.
Functionality F.2 $\mathcal{F}_{\mathsf{mul}}$ Input: The functionality receives inputs $[\![\,a\,]\!]$ and $[\![\,b\,]\!]$. Output: Compute the following 1. Reconstruct $a$ and $b$ to compute $a\cdot b$. 2. Distribute shares $[\![\,ab\,]\!]$ to the parties.
Functionality F.3 $\mathcal{F}_{\mathsf{trunc}}$ Input: The functionality receives input $[\![\,x\,]\!]$. Output: Compute the following 1. Reconstruct $x$ and truncate the lower $f$ bits to get $x^{\prime}$. 2. Distribute shares $[\![\,x^{\prime}\,]\!]$ to the parties.
@@
Functionality F.5 $\mathcal{F}_{\mathsf{gradient}}$ Input: The functionality receives input samples $[\![\,D\,]\!]$, model parameters $[\![\,\theta\,]\!]$, and projection matrix $[\![\,G\,]\!]$. Output: Compute the following 1. Reconstruct $D$, $\theta$, and $G$, and compute the (averaged) gradient $\nabla\theta$ on $D$ for model parameters $\theta$. 2. Compute the lower-dimensional projection $g\leftarrow G\cdot\nabla\theta$. 3. Distribute shares $[\![\,g\,]\!]$ to the parties.
Functionality F.6 $\mathcal{F}_{\mathsf{fp-rsqrt}}$ Input: The functionality receives input $[\![\,x\,]\!]$. Output: Compute the following. 1. Reconstruct $x$ and compute $x^{\prime}\leftarrow 1/{\sqrt{x}}$. 2. Distribute shares $[\![\,x^{\prime}\,]\!]$ to the parties.
@@ Appendix C Additional Experiment Details
CIFAR-10 ResNet-18 Scratch 30 64 $2\times 10^{-4}$ 0.01 0.9 $0.1@[10,20,25]$
Fashion MNIST ConvNet Scratch 30 64 $2\times 10^{-4}$ 0.001 0.9 $0.1@[10,20,25]$
@@ C.1 Detailed Training Configurations
We report detailed training configurations in Table 5. Traceback configurations for each learning setting are given in Table 6.
@@ D.1 Ablations
We evaluate the sensitivity of our tool to changes in the hyperparameter $k$ and the total number of users $m$. For simplicity, we consider only the BadNets and Subpopulation attacks against CIFAR-10 using the same training and attack configurations as in Section 5.2.
Hyperparameter $k$ Ablation. We conduct an ablation on UTrace hyperparameter $k$. To remove the impact of the inner-product heuristic at different values of $k$, we use the simpler $\mathrm{top\text{-}}k$ influence metric. Results are shown in Figure 6. Different attacks have different responses to changes in $k$. We observe that BadNets experiences a stark drop in ranking performance with $k\geq 64$, whereas ranking of users under Subpopulation attacks is more stable with large $k$. A similar observation holds for TPR at low FPR: for BadNets, TPR drops from nearly 1.0 at $k=32$ to under 0.4 at $k=64$ with 4 poisoned datasets. However, the TPR for subpopulation attacks with 4 poisoned users is maximized at the larger $k=256$.
@@
20 Users Ablation. We evaluate the dependence of UTrace on the total number of users by simulating a scenario with 20 total users. We consider scenarios with 2, 4, 6, and 8 poisoned users. Results are given in Tables 7 and 8. We notice slight degradation in ranking performance for BadNets with our method, whereas unlearning maintains perfect rankings. On the other hand, using thresholds determined from FPR analysis produces relatively large FPR for unlearning, qualitatively similar to findings with 10 users. For Subpopulation attacks, the degradation under unlearning is more severe than with 10 users: a mAP of 0.903 with 4 poisoned owners and 20 users compared with 0.976 with 10 users. Our method, on the other hand, actually improves in performance with 20 users, maintaining $\geq 0.991$ mAP with up to 8 poisoned users. This trend is also reflected in the classification metrics: for Subpopulation attacks, unlearning TPR drops to 0 with 8 poisoned datasets, whereas our method achieves a TPR of 0.6022, which is better than with only 10 total users.
@@ D.2 Fashion MNIST
We report the scores for the Fashion-MNIST / ConvNet training scenario in Table 9. In this setting, unlearning outperforms the gradient-based score, with perfect orderings across all attacks. Our method still achieves $\geq 0.943$ mAP and $\geq 0.943$ mRR.
We report results for Fashion-MNIST in Table 10. Similarly to Table 9, we observe a drop in detection rate on BadNets, achieving 72.5% TPR with 4 poisoned datasets. Unlearning achieves a FPR of over 0.8% on BadNets with 1 poisoned dataset, higher than the desired 0.1% for which the threshold was tuned, qualitatively matching the observation on CIFAR-10 in Section 5.2 that tuning thresholds for unlearning’s loss-based score can be difficult.
@@ D.3 Text and Malware Data Modalities
Results. We report the ranking metrics for attacks against SST-2 / $\text{RoBERTa}_{\textsc{base}}$ in Table 16 and the classification metrics in Table 13. We report the ranking metrics for attacks against Ember / EmberNN in Table 15 and the classification metrics in Table 14. We also plot attack-conditioned ROC curves in Figures 9 and 10.
@@
Despite the weaker ranking metrics under the BadNets attack with one poisoned dataset, our method still achieves better a higher TPR of 0.3281 compared to 0.0312. This discrepancy echoes the findings with the BadNets attack in Section 5.2, namely that the thresholds for the loss-based score used by approximate unlearning can depend heavily on the particular attack being considered.

### v2

@@ A.2 True Unlearning Failure Modes
Given the deficiencies of approximate unlearning demonstrated in Section˜A.1, it may seem tempting to replace it with an unlearning mechanism which more accurately approximates leave-one-out retraining. However, we show that in fact true unlearning exacerbates the ability of the adversary to hide poisons among multiple poisoned parties.
@@
In the case when the auditor produces the unlearned models $\theta_{i}^{\prime}$ by retraining from scratch for each user $i$, each time leaving out the user dataset $D_{i}$, the effect of duplicated poisons across multiple user datasets is exacerbated. We consider a minimal theoretical example demonstrating this possibility. Consider two distributions $\mu_{c}$ and $\mu_{p}$ over $\mathcal{X}\times\mathcal{Y}$, representing the clean and poisoned distributions, respectively. If the distributions $\mu_{c}$ and $\mu_{p}$ are compatible with each other, and if the model space $\Theta$ is sufficiently rich to simultaneously learn both distributions, then learning on any nontrivial mixture of those distributions yields the same set of loss-minimizing classifiers. Intuition for this idea is given Figure˜5. We formalize this idea below:
@@
Implications of Theorem˜A.2. The distributions $\mu_{c}$ and $\mu_{p}$ in Theorem˜A.2 can be replaced with the empirical distributions $D_{i}$ for each user. Interpreted this way, the theorem describes a simplified setting where each user dataset is one of two possible datasets, $D_{c}$ or $D_{p}$. If the necessary assumptions are satisfied, performing true unlearning on any individual dataset that is duplicated across multiple users corresponds to setting a different $\alpha\in(0,1)$ (namely, $\alpha=\frac{a-1}{a+b-1}$, where $a$ and $b$ are the multiplicities of each dataset and we have removed one from those counted by $a$).
@@ B.1 Ideal Functionalities
[…]
@@
[…]
@@ Appendix C Additional Experiment Details
CIFAR-10 ResNet-18 Scratch 30 64 $2\times 10^{-4}$ 0.01 0.9 0.1@[10,20,25]0.1@[10,20,25]0.1 @ [ 10 , 20 , 25 ]
Fashion MNIST ConvNet Scratch 30 64 $2\times 10^{-4}$ 0.001 0.9 0.1@[10,20,25]0.1@[10,20,25]0.1 @ [ 10 , 20 , 25 ]
@@ C.1 Detailed Training Configurations
We report detailed training configurations in Table˜5. Traceback configurations for each learning setting are given in Table˜6.
@@ D.1 Ablations
We evaluate the sensitivity of our tool to changes in the hyperparameter $k$ and the total number of users $m$. For simplicity, we consider only the BadNets and Subpopulation attacks against CIFAR-10 using the same training and attack configurations as in Section˜5.2.
Hyperparameter $k$ Ablation. We conduct an ablation on UTrace hyperparameter $k$. To remove the impact of the inner-product heuristic at different values of $k$, we use the simpler $\mathrm{top\text{-}}k$ influence metric. Results are shown in Figure˜6. Different attacks have different responses to changes in $k$. We observe that BadNets experiences a stark drop in ranking performance with $k\geq 64$, whereas ranking of users under Subpopulation attacks is more stable with large $k$. A similar observation holds for TPR at low FPR: for BadNets, TPR drops from nearly 1.0 at $k=32$ to under 0.4 at $k=64$ with 4 poisoned datasets. However, the TPR for subpopulation attacks with 4 poisoned users is maximized at the larger $k=256$.
@@
20 Users Ablation. We evaluate the dependence of UTrace on the total number of users by simulating a scenario with 20 total users. We consider scenarios with 2, 4, 6, and 8 poisoned users. Results are given in Tables˜7 and 8. We notice slight degradation in ranking performance for BadNets with our method, whereas unlearning maintains perfect rankings. On the other hand, using thresholds determined from FPR analysis produces relatively large FPR for unlearning, qualitatively similar to findings with 10 users. For Subpopulation attacks, the degradation under unlearning is more severe than with 10 users: a mAP of 0.903 with 4 poisoned owners and 20 users compared with 0.976 with 10 users. Our method, on the other hand, actually improves in performance with 20 users, maintaining 0.991$\geq$ 0.991≥ 0.991 mAP with up to 8 poisoned users. This trend is also reflected in the classification metrics: for Subpopulation attacks, unlearning TPR drops to 0 with 8 poisoned datasets, whereas our method achieves a TPR of 0.6022, which is better than with only 10 total users.
@@ D.2 Fashion MNIST
We report the scores for the Fashion-MNIST / ConvNet training scenario in Table˜9. In this setting, unlearning outperforms the gradient-based score, with perfect orderings across all attacks. Our method still achieves 0.943$\geq$ 0.943≥ 0.943 mAP and 0.943$\geq$ 0.943≥ 0.943 mRR.
We report results for Fashion-MNIST in Table˜10. Similarly to Table˜9, we observe a drop in detection rate on BadNets, achieving 72.5% TPR with 4 poisoned datasets. Unlearning achieves a FPR of over 0.8% on BadNets with 1 poisoned dataset, higher than the desired 0.1% for which the threshold was tuned, qualitatively matching the observation on CIFAR-10 in Section˜5.2 that tuning thresholds for unlearning’s loss-based score can be difficult.
@@ D.3 Text and Malware Data Modalities
Results. We report the ranking metrics for attacks against SST-2 / $\text{RoBERTa}_{\textsc{base}}$ in Table˜16 and the classification metrics in Table˜13. We report the ranking metrics for attacks against Ember / EmberNN in Table˜15 and the classification metrics in Table˜14. We also plot attack-conditioned ROC curves in Figures˜9 and 10.
@@
Despite the weaker ranking metrics under the BadNets attack with one poisoned dataset, our method still achieves better a higher TPR of 0.3281 compared to 0.0312. This discrepancy echoes the findings with the BadNets attack in Section˜5.2, namely that the thresholds for the loss-based score used by approximate unlearning can depend heavily on the particular attack being considered.